OpenStudy (anonymous):

please help!!!!!! Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6.

jimthompson5910 (jim_thompson5910):

the slope of y = -x + 6 is what?

OpenStudy (anonymous):

is it -x

jimthompson5910 (jim_thompson5910):

it's just a number (not a variable)

jimthompson5910 (jim_thompson5910):

it's the number out front of the x -x is the same as -1x

jimthompson5910 (jim_thompson5910):

so the slope is -1

jimthompson5910 (jim_thompson5910):

what is the perpendicular slope?

OpenStudy (anonymous):

-4,-2

jimthompson5910 (jim_thompson5910):

-1 = -1/1 flip the fraction and flip the sign to get 1/1 = 1

jimthompson5910 (jim_thompson5910):

so the perpendicular slope is 1

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):

use this, along with the given point, to find the equation of the perpendicular line

OpenStudy (anonymous):

so do i have to use the slope intercept form

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

would it be y = x - 6

jimthompson5910 (jim_thompson5910):

y = mx + b y = 1x + b -2 = 1(-4) + b solve for b

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so it would be y = x + 2

jimthompson5910 (jim_thompson5910):

yep, you got it

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np